Hi again,
I could also use some help with this problem:

Find the equation of the line that passes through the point A(-2, 5/6) with slope -7/6. Express the answer in general form.

Thanks again,
Lou

Answer by checkley71(8403)

Y=mX+b or 5/6=-7/6*-2+b or 5/6=14/6+b or 5/6-14/6=b or-9/6=b or b=-1.5 thus
y=-7/6x-1.5